OpenStudy (anonymous):

Please help I don't understand this at all..... A right prism with a surface area of 82 square feet and a volume of 42 cubic feet; scale factor 1:2

OpenStudy (anonymous):

Is that the entirety of the question?

OpenStudy (amistre64):

is a right prism just a cube?

OpenStudy (amistre64):

not if it scales to 1:2....got it ;)

OpenStudy (anonymous):

The solid described is similar to a large solid with the given scale factor. Find the surface area and volume of a larger solid

OpenStudy (amistre64):

the shape of the solid doesnt matter? or does it. a right prism can be anything as long as it is the same shape stacked up on top of itself.... or am I mistaken?

OpenStudy (amistre64):

the easies thing to do is to make it a rectangular prism...i think

OpenStudy (amistre64):

82/6 would be a square prism with 6 equal sides ...

OpenStudy (amistre64):

but that doesnt work out evenly

OpenStudy (amistre64):

Volume = l*w*h = 42 As = Area of faces

OpenStudy (amistre64):

im at a loss....sorry...

OpenStudy (amistre64):

maybe; 4 side + 2 bases = 82 that would still be a rectanule prism,,,,

OpenStudy (amistre64):

if it had square bases then 4wh + 2w = 82 would be simpler

OpenStudy (amistre64):

4wh + 2w^2 = 82 42 = w^2h h = 42/w^2 4(42/w) + 2w^2 = 82

OpenStudy (amistre64):

168/w + 2w^2 = 82 168 + 2w^3 = 82w 2w^3 -82w +168 = 0 2(w^3 - 41w + 84) = 0 w^3 - 41w + 84 = 0 is what I got so far if I did it right

OpenStudy (amistre64):

this just seems harder than what it should be called for :)

OpenStudy (amistre64):

and it looks like my w is going negative....... which doesnt bode well for my plans.....

OpenStudy (amistre64):

if we knew the original shape for the base of this right prism, id be better to calculate..... other than that....sorry